# A Case for Library-Level k-Means Binning in Histogram Gradient-Boosted Trees

### v1

@@ 3.2 Synthetic diagnostics
All synthetic experiments were executed on a MacBook Pro (Apple M1 Pro, 8 cores, 16 GB RAM, macOS 15.0). They completed in about 1.55 hours of wall-clock time, used roughly 4.6 core-hours in total, and never exceeded 0.5 GB of memory. Preliminary experiments roughly tripled this compute cost. These compute costs are essentially negligible in comparison to the much larger compute requirements of testing on real-world benchmarks.
@@
Conversely, Fig. 2b fixes the sample size while varying the number of bins. Here, the binning methods differ even without outliers; k-means outperforms quantile on a pure normally-distributed dataset by 43% when using only 16 bins. With significant outliers, we see similar results as in experiment (4); k-means outperforms quantile by nearly 90% when outliers constitute 1-5% with 16-32 bins. The top row (100% “outliers") corresponds to a data set drawn entirely from an exponential distribution. Despite its heavy right-hand tail, it behaves similar to the pure-Gaussian case—especially once the bin budget exceeds about 64—indicating that k-means’ advantage is triggered by the presence of a small, label-relevant tail alongside a dense bulk, not by heavy-tailedness per se.
@@ 4 Computational Efficiency
Figure 3 benchmarks the bin-construction pass on a single-core Apple M1 for uniformly distributed data ranging from to rows. Equal-width and quantile curves remain close across the range, suggesting essentially equivalent big-O time complexity and constant factor. -means with quantile seeding follows the same slope but with a higher intercept: on ten million rows quantile finishes in 2.5s, whereas -means requires 5.1s; at one million rows the gap is a mere 0.3s. We considered testing k-means binning with k-means++ initialization, but found a significant increase in time complexity with no performance improvements. Because the cost to apply k-means binning with quantile initialization is incurred once, can be cached, and is typically dwarfed by cross-validation or hyper-parameter sweeps, the extra 3 seconds on large datasets are negligible in practice—especially relative to the strong error reduction we observe on skewed regression tasks.
@@ 5 Conclusion
This work revisited a seemingly innocuous design choice that underlies all modern histogram-based GBDT implementations: how should one place the bin boundaries? Our systematic study across 33 real-world tasks and a five-factor synthetic suite shows that the long-standing default of equal-frequency (quantile) cuts is not sacrosanct.
We find that, in datasets with small, label-relevant tails, or in scenarios with low bin budgets, k-means can cut error significantly, sometimes by more than half. In particular, we find that k-means matches exact splitting in benign cases and consistently outperforms quantile whenever tails or tight bin budgets make histograms too coarse.
Because training time after binning is identical for all binning methods, the only extra cost is a one-off, cacheable preprocessing pass—essentially negligible at roughly 2s for 10 M rows on a single Apple M1 core.
@@ 6 Limitations and Future Work
We deliberately ran our study with scikit-learn’s "vanilla" histogram GBDT, eliminating all library-specific preprocessing. A small replica (found in Appendix 5) using XGBoost’s exact method shows the same pattern, but a full integration into production pipelines—e.g. LightGBM’s exclusive-feature bundling, XGBoost’s quantile sketch, CatBoost’s ordered targets—remains an engineering task for future work. Our results therefore speak most directly to library maintainers, who can offer a bin_method=k-means flag once these code paths are updated.
As shown throughout, k-means outperforms quantile most when datasets have a small, label-relevant tail or when bin budget is tight. If the tail information can already be reconstructed from other features, or if the target changes sharply at unknown boundaries inside a dense bulk (e.g. a normally distributed predictor with a step change at ), the advantage shrinks considerably. Likewise, if users have access to large bin budgets, they should expect to see smaller gaps between k-means and quantile binning. Finally, our research is primarily empirical; we leave a full theoretical characterization of the heavy-tail advantage to future work.
Appendix A Additional Experimental Details
A.1 Dataset Description
The following table describes observations and features from each dataset. These datasets range from observations, and up to 400+ features.
| Dataset | #Obs | #Feat |
| cpu_act | 8192 | 22 |
| pol | 15000 | 27 |
| elevators | 16599 | 17 |
| wine_quality | 6497 | 12 |
| Ailerons | 13750 | 34 |
| houses | 20640 | 9 |
| house_16H | 22784 | 17 |
| diamonds | 53940 | 7 |
| Brazilian_houses | 10692 | 9 |
| Bike_Sharing_Demand | 17379 | 7 |
| nyc-taxi-green-dec-2016 | 581835 | 10 |
| house_sales | 21613 | 16 |
| sulfur | 10081 | 7 |
| medical_charges | 163065 | 4 |
| MiamiHousing2016 | 13932 | 14 |
| superconduct | 21263 | 80 |
| yprop_4_1 | 8885 | 43 |
| abalone | 4177 | 8 |
| delays_zurich_transport | 5465575 | 9 |
| Dataset | #Obs | #Feat |
| credit | 16714 | 11 |
| electricity | 38474 | 8 |
| covertype | 566602 | 11 |
| pol | 10082 | 27 |
| house_16H | 13488 | 17 |
| MagicTelescope | 13376 | 11 |
| bank-marketing | 10578 | 8 |
| MiniBooNE | 72998 | 51 |
| Higgs | 940160 | 25 |
| eye_movements | 7608 | 21 |
| Diabetes130US | 71090 | 8 |
| jannis | 57580 | 55 |
| default-of-credit-card-clients | 13272 | 21 |
| Bioresponse | 3434 | 420 |
| california | 20634 | 9 |
| heloc | 10000 | 23 |
A.2 Hyper-parameter Search Space
@@
A.3 Real-world Benchmark with Low Bin Budget
To understand the effect of bin budget on real-world datasets, we conducted equivalent experiments as described in section 3.1 with . We chose this value because it is the recommended bin size for training GBDTs on a GPU as described by Zhang et al. (2017). Note specifically the new gaps in the CPU Act and Pol datasets (10% and 2.7% respectively), and the even larger gap on Brazilian Houses (68%). Quantile now wins on two datasets, by .
| Dataset Name | Quantile | Uniform | k-means | Exact |
@@ A.3 Real-world Benchmark with Low Bin Budget
A.4 Re-running Representative Datasets on XGBoost
To confirm that our experiments carry over to commercial GBDT algorithms, we compared our k-means binner with quantile and uniform on three representative regression datasets. We find that k-means ties once again on CPU Act and Superconduct, while achieving even greater wins on Brazilian Houses.
| Dataset Name | Quantile | Uniform | k-means | Exact |
@@ A.4 Re-running Representative Datasets on XGBoost
A.5 Synthetic-Data Generator
Algorithm 1 shows the procedure used to create the five diagnostic suites discussed in section 3.2. We set , so adjacent Gaussian modes overlap by roughly of their mass—enough to avoid empty bins yet still preserve distinct peaks.

### v2

@@ 4.2 Synthetic diagnostics
All synthetic experiments were executed on a MacBook Pro (Apple M1 Pro, 8 threads, 16 GB RAM, macOS 15.0). They completed in about 1.55 hours of wall-clock time, used roughly 4.6 core-hours in total, and never exceeded 0.5 GB of memory. Preliminary experiments roughly tripled this compute cost. These compute costs are essentially negligible in comparison to the much larger compute requirements of testing on real-world benchmarks.
@@
Conversely, Fig. 2b fixes the sample size while varying the number of bins. Here, the binning methods differ even without outliers; k-means outperforms quantile on a pure normally-distributed dataset by 43% when using only 16 bins. With significant outliers, we see similar results as in experiment (4); k-means outperforms quantile by nearly 90% when outliers constitute 1-5% with 16-32 bins. The top row corresponds to a data set drawn entirely from an exponential distribution. Despite its heavy right-hand tail, it behaves similar to the pure-Gaussian case—especially once the bin budget exceeds about 64—indicating that k-means’ advantage is triggered by the presence of a small, label-relevant tail alongside a dense bulk, not by heavy-tailedness per se.
@@ 5 Computational Efficiency
Figure 3 benchmarks the bin-construction pass on a single-thread Apple M1 for uniformly distributed data ranging from to rows. Times are per feature; to estimate wall-clock cost for columns, multiply the reported value by , as both 1-D binning methods and GBDTs scale linearly in . Equal-width and quantile curves remain close across the range, suggesting essentially equivalent big-O time complexity and constant factor. -means with quantile seeding follows the same slope but with a higher intercept: on ten million rows quantile finishes in 2s, whereas -means requires 5.5s; at one million rows the gap is a mere 0.3s. We also experimented with -means++ seeding (Arthur & Vassilvitskii, 2007), but saw no meaningful accuracy improvement over -means with quantile seeding. Since the extra initialization pass in -means++ made preprocessing substantially slower, we decided to use quantile seeding throughout.
We note that quantile, uniform, and quantile-seeded -means consume less than one-third of total training time. This will remain the case regardless of the number of features , since both binning and GBDT training itself scale linearly in . In addition, as the cost to bin is incurred once and can be cached, the extra 3.5 seconds per feature on 10M+ row datasets are negligible in practice—especially relative to the strong error reduction we observe on skewed regression tasks. For datasets that exceed available RAM, a streaming 1-D -means variant (Sculley, 2010) can be substituted without changing the API; we leave evaluating this option on truly massive tables to future work.
@@ 6 Conclusion
This work revisited a seemingly innocuous design choice that underlies all modern histogram-based GBDT implementations: how should one place the bin boundaries? Our systematic study across 33 real-world tasks, a five-factor synthetic suite, and a new lower-bound theorem, shows that the long-standing default of equal-frequency cuts is not sacrosanct.
We find that, in datasets with small, label-relevant tails, or in scenarios with low bin budgets, k-means can cut error significantly, sometimes by more than half. In particular, we find that k-means matches exhaustive splitting in benign cases and consistently outperforms quantile whenever tails or tight bin budgets make histograms too coarse.
Because training time after binning is identical for all binning methods, the only extra cost is a one-off, cacheable preprocessing pass—essentially negligible at roughly 3.5s for 10M rows on a single Apple M1 thread.
@@ 7 Limitations and Future Work
We deliberately ran our study with scikit-learn’s "vanilla" histogram GBDT, eliminating all library-specific preprocessing. A small replica (found in Appendix F) using XGBoost’s exact method shows the same pattern, but a full integration into production pipelines—e.g. LightGBM’s exclusive-feature bundling, XGBoost’s quantile sketch, CatBoost’s ordered targets—remains an engineering task for future work. Our results therefore speak most directly to library maintainers, who can offer a bin_method=k-means flag once these code paths are updated.
As shown throughout, k-means outperforms quantile most when datasets have a small, label-relevant tail or when bin budget is tight. If the tail information can already be reconstructed from other features, or if the target changes sharply at unknown boundaries inside a dense bulk (e.g. a normally distributed predictor with a step change at ), the advantage shrinks considerably. Likewise, if users have access to large bin budgets, they should expect to see smaller gaps between k-means and quantile binning.
Appendix A Proof of Strictness for Linear Model
We begin the proof as in Section 3, with the additional assumption that . In particular, the proof remains the same up until (12), wherein it continues as follows:
Consequently,
Since k-means binning minimizes , it also maximizes the explained variance of Y obtained when treating all values in a given bin as equivalent.
Appendix B Dataset Description
The following table describes observations and features from each dataset. For a given dataset, we calculate the skew by averaging the skew of each of the X columns. These datasets range from observations, and up to 400+ features. We calculate skew using the default scipy Fisher–Pearson sample skewness (scipy.stats.skew, bias=True). Recomputing with the unbiased variant (bias=False) changes values by , leaving all printed numbers unchanged.
| Dataset | Skew | #Obs | #Feat |
| cpu_act | 5.40 | 8 192 | 22 |
| pol | 4.78 | 15 000 | 27 |
| elevators | -0.60 | 16 599 | 17 |
| wine_quality | 1.36 | 6 497 | 12 |
| Ailerons | 0.60 | 13 750 | 34 |
| houses | 2.23 | 20 640 | 9 |
| house_16H | 6.14 | 22 784 | 17 |
| diamonds | 1.03 | 53 940 | 7 |
| Brazilian_houses | 30.23 | 10 692 | 9 |
| Bike_Sharing_Demand | 0.06 | 17 379 | 7 |
| nyc-taxi-green-2016 | 2.56 | 581 835 | 10 |
| house_sales | 2.40 | 21 613 | 16 |
| sulfur | -0.03 | 10 081 | 7 |
| medical_charges | 5.02 | 163 065 | 4 |
| MiamiHousing2016 | 0.93 | 13 932 | 14 |
| superconduct | 0.67 | 21 263 | 80 |
| yprop_4_1 | -0.29 | 8 885 | 43 |
| abalone | 0.62 | 4 177 | 8 |
| zurich_transport | 1.00 | 5 465 575 | 9 |
| Dataset | Skew | #Obs | #Feat |
| credit | 20.17 | 16 714 | 11 |
| electricity | 12.07 | 38 474 | 8 |
| covertype | 0.28 | 566 602 | 11 |
| pol | 5.34 | 10 082 | 27 |
| house_16H | 5.78 | 13 488 | 17 |
| MagicTelescope | 0.58 | 13 376 | 11 |
| bank-marketing | 3.65 | 10 578 | 8 |
| MiniBooNE | -11.68 | 72 998 | 51 |
| Higgs | 1.56 | 940 160 | 25 |
| eye_movements | 3.18 | 7 608 | 21 |
| Diabetes130US | 5.48 | 71 090 | 8 |
| jannis | -0.22 | 57 580 | 55 |
| default-credit-card | 5.55 | 13 272 | 21 |
| Bioresponse | 6.95 | 3 434 | 420 |
| california | 19.56 | 20 634 | 9 |
| heloc | -0.32 | 10 000 | 23 |
Appendix C Hyper-parameter Search Space
@@
Appendix D Small Benchmark for Alternative Binners
Although the body of the paper focuses on quantile, uniform and Lloyd–-means, we ran a miniature benchmark to gauge the price of more optimal schemes. In particular, we compared
Given the extended training time of 1-D -means, we restricted the test to a small (Abalone, ) and a medium (Diamonds, ) dataset. All runs use the hyper-parameter grid in Appendix C. Binning time is reported per split; the full CV grid therefore costs the numbers shown.
| Binning Method | Abalone (0.22s to train) | Diamonds (1.82s to train) | ||
| Binning time (s) | MSE | Binning time (s) | MSE | |
| Quantile | 0.003 | 4.76 | 0.03 | 5.30 |
| -means | 0.028 | 4.65 | 0.39 | 5.36 |
| MILP | 0.161 | 5.42 | 0.19 | 8.80 |
| 1-D -means optimal | 0.559 | 4.73 | 6.75 | 5.31 |
Take-away.
MILP-optimal runs in a reasonable time but often merges features down to bins, hurting accuracy; DP -means matches Lloyd on error but is 15x slower (and >2x slower than training itself). Therefore, neither option is competitive for large-scale histogram GBDT training.
Appendix E Real-world Benchmark with Low Bin Budget
To understand the effect of bin budget on real-world datasets, we conducted equivalent experiments as described in section 4.1 with . In LightGBM’s GPU implementation, Zhang et al. (2017, p.7, p.13) note that using fewer bins (6-bit packing with one reserved value, i.e., 63 usable bins) can increase available per-workgroup workload and reduce local memory requirements, both of which speed up training. They report significant speedups at this setting, so we rerun our real-world regression suite with B=63 to reflect the GPU-relevant regime and report those results below.
Note specifically the new gaps in the CPU Act and Pol datasets (10% and 2.7% respectively), and the even larger gap on Brazilian Houses (68%). Quantile now wins on two datasets, by . We hypothesize two mechanisms for quantile’s two small wins in the low-bin regime. First, k-means can allocate the same number of bins to extreme values as at higher B, but when B is small those bins consume a much larger proportion of the total budget – which can sometimes be counterproductive if the tails are not label-relevant. Second, quantile binning may simply win by chance – in low-bin regimes, model performance becomes sensitive to dataset quirks, so accidental alignments between bin edges and label structure can drive small empirical wins.
| Dataset Name | Quantile | Uniform | k-means | Exhaustive |
@@ Appendix E Real-world Benchmark with Low Bin Budget
Appendix F Re-running Representative Datasets on XGBoost
To confirm that our experiments carry over to commercial GBDT algorithms, we compared our k-means binner with quantile and uniform on three representative regression datasets. We find that -means achieves a greater win on Brazilian Houses, while remaining statistically indistinguishable from quantile () on CPU Act and Superconduct.
| Dataset Name | Quantile | Uniform | k-means | Exhaustive |
@@ Appendix F Re-running Representative Datasets on XGBoost
Appendix G Synthetic-Data Generator
Algorithm 1 shows the procedure used to create the five diagnostic suites discussed in Section 4.2. We motivated this algorithm as a combination of the simplest building blocks that yield the distributional features we wanted to probe (multi-modality and label-relevant outliers).
First, a Gaussian mixture with n_modes components creates a controllable number of modes. We fix dist=4 so that adjacent modes share of their mass – enough to avoid empty bins while keeping the peaks distinct.
Second, we standardize each feature to equalize scale before any outliers are introduced.
Third, an exponential distribution of size applied to a p_out fraction of samples generates controllable outlier magnitude and frequency.
Finally, a simple linear target () ensures the exponential tails are label-relevant (i.e. extreme maps to extreme ).
Appendix H Variability of Experiments
Here we provide copies of all previous experimental tables in the main text as well as in the appendix, with standard errors of the mean across random states included. As before, bolded values imply statistical significance over the second-best method at p = 0.05, after application of Benjamini–Hochberg.
H.1 255-bin results
| Dataset Name | Quantile | Uniform | k-means | Exhaustive |
| Regression (MSE) | ||||
| cpu_act | 5.043 (0.170) | 5.094 (0.105) | 4.965 (0.107) | 5.092 (0.122) |
| pol | 3.337 (0.086) | 3.337 (0.086) | 3.337 (0.086) | 3.347 (0.088) |
| elevators | 4.861 (0.067) | 4.842 (0.063) | 4.862 (0.068) | 4.881 (0.068) |
| wine_quality | 4.096 (0.050) | 4.089 (0.046) | 4.128 (0.054) | 4.117 (0.053) |
| Ailerons | 2.518 (0.022) | 2.526 (0.021) | 2.523 (0.020) | 2.519 (0.021) |
| houses | 5.274 (0.047) | 5.358 (0.046) | 5.283 (0.046) | 5.292 (0.040) |
| house_16H | 3.338 (0.111) | 3.674 (0.133) | 3.415 (0.136) | 3.262 (0.132) |
| diamonds | 5.458 (0.020) | 5.602 (0.021) | 5.464 (0.022) | 5.459 (0.022) |
| Brazilian_houses | 5.382 (1.087) | 20.272 (0.520) | 2.433 (0.625) | 2.156 (0.576) |
| Bike_Sharing_Demand | 9.697 (0.041) | 9.697 (0.041) | 9.697 (0.041) | 9.694 (0.042) |
| nyc-taxi-green-dec-2016 | 1.553 (0.011) | 1.959 (0.008) | 1.522 (0.010) | 1.320 (0.016) |
| house_sales | 3.183 (0.021) | 3.217 (0.022) | 3.175 (0.022) | 3.206 (0.020) |
| sulfur | 4.649 (0.361) | 4.698 (0.407) | 4.663 (0.358) | 4.779 (0.383) |
| medical_charges | 6.686 (0.035) | 7.153 (0.042) | 6.600 (0.036) | 6.584 (0.035) |
| MiamiHousing2016 | 2.259 (0.032) | 2.291 (0.034) | 2.264 (0.032) | 2.309 (0.031) |
| superconduct | 1.012 (0.013) | 1.036 (0.014) | 1.015 (0.014) | 1.038 (0.013) |
| yprop_4_1 | 9.457 (0.530) | 9.481 (0.534) | 9.464 (0.532) | 9.474 (0.527) |
| abalone | 4.775 (0.066) | 4.788 (0.062) | 4.767 (0.068) | 4.759 (0.068) |
| Regression MRR | 0.72 | 0.43 | 0.65 | |
| Classification (ROC AUC) | ||||
| credit | 0.857 (0.001) | 0.827 (0.001) | 0.857 (0.001) | 0.857 (0.001) |
| electricity | 0.950 (0.001) | 0.918 (0.001) | 0.948 (0.001) | 0.960 (0.001) |
| covertype | 0.933 (0.003) | 0.931 (0.003) | 0.932 (0.003) | 0.931 (0.003) |
| pol | 0.999 (0.000†) | 0.999 (0.000†) | 0.999 (0.000†) | 0.999 (0.000†) |
| house_16H | 0.951 (0.001) | 0.947 (0.001) | 0.951 (0.001) | 0.950 (0.001) |
| MagicTelescope | 0.931 (0.001) | 0.931 (0.001) | 0.931 (0.001) | 0.930 (0.001) |
| bank-marketing | 0.886 (0.001) | 0.886 (0.001) | 0.886 (0.001) | 0.886 (0.001) |
| MiniBooNE | 0.983 (0.000†) | 0.967 (0.000†) | 0.983 (0.000†) | 0.982 (0.000†) |
| eye_movements | 0.705 (0.003) | 0.709 (0.003) | 0.709 (0.003) | 0.718 (0.003) |
| Diabetes130US | 0.647 (0.001) | 0.647 (0.001) | 0.647 (0.001) | 0.647 (0.001) |
| jannis | 0.868 (0.001) | 0.867 (0.001) | 0.868 (0.001) | 0.867 (0.001) |
| default-of-credit-card-clients | 0.781 (0.002) | 0.779 (0.002) | 0.781 (0.002) | 0.780 (0.002) |
| Bioresponse | 0.861 (0.002) | 0.859 (0.002) | 0.860 (0.002) | 0.858 (0.002) |
| california | 0.967 (0.000†) | 0.962 (0.001) | 0.967 (0.000†) | 0.966 (0.000†) |
| heloc | 0.798 (0.002) | 0.798 (0.002) | 0.798 (0.002) | 0.798 (0.002) |
| Classification MRR | 0.68 | 0.41 | 0.70 | |
H.2 63-bin results
| Dataset Name | Quantile | Uniform | k-means | Exhaustive |
| Regression (MSE) | ||||
| cpu_act | 5.579 (0.203) | 5.416 (0.145) | 5.140 (0.126) | 5.092 (0.122) |
| pol | 3.483 (0.087) | 3.556 (0.087) | 3.421 (0.084) | 3.347 (0.088) |
| elevators | 4.936 (0.060) | 4.927 (0.061) | 4.859 (0.062) | 4.881 (0.068) |
| wine_quality | 4.142 (0.054) | 4.168 (0.049) | 4.114 (0.051) | 4.117 (0.053) |
| Ailerons | 2.529 (0.021) | 2.525 (0.021) | 2.510 (0.019) | 2.519 (0.021) |
| houses | 5.451 (0.042) | 6.030 (0.046) | 5.409 (0.046) | 5.292 (0.040) |
| house_16H | 3.419 (0.135) | 3.881 (0.130) | 3.398 (0.126) | 3.262 (0.132) |
| diamonds | 5.501 (0.021) | 5.810 (0.019) | 5.526 (0.021) | 5.459 (0.022) |
| Brazilian_houses | 6.507 (1.208) | 34.465 (0.850) | 2.222 (0.640) | 2.156 (0.576) |
| Bike_Sharing_Demand | 9.694 (0.046) | 9.693 (0.046) | 9.675 (0.044) | 9.694 (0.042) |
| nyc-taxi-green-dec-2016 | 1.767 (0.011) | 2.268 (0.007) | 1.763 (0.013) | 1.320 (0.016) |
| house_sales | 3.195 (0.024) | 3.322 (0.023) | 3.189 (0.022) | 3.206 (0.020) |
| sulfur | 5.089 (0.379) | 4.627 (0.389) | 4.589 (0.349) | 4.779 (0.383) |
| medical_charges | 7.365 (0.032) | 13.980 (0.126) | 6.737 (0.037) | 6.584 (0.035) |
| MiamiHousing2016 | 2.267 (0.031) | 2.312 (0.033) | 2.289 (0.029) | 2.309 (0.031) |
| superconduct | 9.931 (0.131) | 10.825 (0.131) | 10.193 (0.139) | 10.383 (0.133) |
| yprop_4_1 | 9.425 (0.535) | 9.432 (0.538) | 9.435 (0.531) | 9.474 (0.527) |
| abalone | 4.775 (0.067) | 4.837 (0.066) | 4.790 (0.065) | 4.759 (0.068) |
| Regression MRR | 0.59 | 0.39 | 0.85 | |
H.3 XGB results
| Dataset Name | Quantile | Uniform | k-means | Exhaustive |
| Regression (MSE) | ||||
| cpu_act | 5.047 (0.145) | 5.204 (0.166) | 5.454 (0.193) | 5.631 (0.229) |
| Brazilian_houses | 4.555 (0.913) | 19.303 (0.445) | 1.655 (0.407) | 1.639 (0.393) |
| superconduct | 1.003 (0.022) | 1.038 (0.022) | 1.014 (0.021) | 1.032 (0.021) |